Two dice are rolled. Find the indicated probability.
What is the probability that the sum is even if the first die is 4?

Answer by checkley71(8403)

4+2=6 EVEN
4+4=8 EVEN
4+6=10 EVEN
ANSWER IS 3/6=.5 OR 50%